For an iron - core transformer with Ep =200 V; Es =2400 V; and the primary coil consists of 50 turns. Find the secondary coils turns:


Select one:

A. 600 turns

B. None of these

C. 10 turns

D. 300 turns


1
Expert's answer
2021-06-11T06:11:10-0400

\frac {E_p} {E_s}= \frac {N_p} {N_s}

E_p = 200V

E_s = 2400V

N_p = 50 turns

N_s = ?

\frac {200} {2400} = \frac {50} {N_s}

N_s = \frac {50 \times 2400} {200}

N_s = 600 turns

Answer

(A)